# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ VIII Conclusion
The semiparametric statistical efficiency in estimation problems for elliptically symmetric distributed data was analyzed in this paper. In particular, we studied the impact of finite and infinite-dimensional nuisance parameters can have on the estimation of the parameters of interest which, in the case of elliptical distributions, are the location vector $\bm{\mu}$ and the covariance matrix $\bm{\Sigma}$. The profound and counter-intuitive result that emerged is that, in the presence of specific finite-dimensional nuisance parameters, semiparametric efficiency can be equivalent to parametric efficiency. Specifically, in the case of elliptical distributions, not knowing the density generator does not cause any loss of efficiency when estimating $\bm{\mu}$ or a scaled version of $\bm{\Sigma}$. This result had already been demonstrated by Hallin and Paindaveine using Le Cam’s asymptotic theory in [9, 10]. Unlike these works, in this article we used a purely geometric approach based on Hilbert spaces. Furthermore, as an advancement over the state of the art, we analyzed the case, important in many applications, in which the parameters of interest and the finite-dimensional nuisance parameters are given by a parameterization of the location vector and of the covariance matrix. A general condition that the parameterization in question must satisfy in order for the semiparametric efficiency to be equal to the parametric efficiency has been derived in this work. This condition therefore allows us to test this property for any particular parameterizations. Two examples were investigated here, including the well-known low-rank parameterization, often arising in many practical signal processing applications. The paper concluded with a section in which the results derived for RES distributions are extended to the case of C-CES and NC-CES distributions. The natural follow-up to this paper will be on the development of semiparametric estimators capable of achieving parametric efficiency. A promising approach for achieving this goal is that of rank-based ($R$-) estimators [46, 47, 40].
@@
    For a given $\mathbf{h}\in\mathcal{H}_{q}$, the projection $\Pi(\mathbf{h}|\mathcal{H}_{q})$ is defined as the vector in $\mathcal{H}_{q}$ such that [16, Theo. 5.1]:
@@
    The $q$-replicating Pythagorean theorem [13, Theo. 3.3]: Let $\mathcal{H}_{q}$ and $\mathcal{U}_{q}\subseteq\mathcal{H}_{q}$ the $q$-replicating Hilbert space and subspace defined as in Lemma .1. Let us take $\mathbf{h}\in\mathcal{H}_{q}$ and $\mathbf{u}\in\mathcal{U}_{q}$. If $\mathbf{h}\perp\mathbf{u}$, then:
@@
    Proof of Lemma II.2. Let us start by noticing that the closed subspaces $(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}$ and $\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}$ are orthogonal. In fact, it is immediate to verify that, each element $h\in\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}$ can be written as $h=h_{2}+h_{3}$ with $h_{2}\in\mathcal{T}_{2}$ and $h_{3}\in\mathcal{T}_{3}$. Moreover, $\forall g\in\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}$ we have that $g\in\mathcal{T}_{2}^{\perp}$ and $g\in\mathcal{T}_{3}^{\perp}$ and consequently $h\perp g$. In addition, $\mathcal{T}_{2}^{\perp}$ can be expressed as the (direct) sum of the and these two closed subspaces, i.e. $\mathcal{T}_{2}^{\perp}=(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}+(\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp})$. Consequently, from the property (20), we get:
@@
    Implicit derivation of $\bm{\nabla}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}$. We follow the procedure discussed in [10, Sect. 4]. Let us start by defining the mapping $v_{11}^{S}:\mathbb{R}^{m(m+1)/2-1}\mapsto\ \mathbb{R}$ implicitly defined by the constraint $S(\mathbf{V}_{S})\equiv S(v_{11}^{S}(\underline{\mathrm{vecs}}(\mathbf{V}_{S})),\underline{\mathrm{vecs}}(\mathbf{V}_{S}))=1$. Thanks to the implicit function theorem, under Assumptions A1, A2, A3, this mapping exists, is unique and continuously differentiable around a given $\mathrm{vecs}(\mathbf{V}_{S})$. Then, we can differentiate both side of $S(v_{11}^{S}(\underline{\mathrm{vecs}}(\mathbf{V}_{S})),\underline{\mathrm{vecs}}(\mathbf{V}_{S}))=1$ w.r.t. $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$ to get:
@@
    The scale $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ implies that $|\mathbf{V}_{S}|=1$. Taking as before the differential on both side, we have [23, pg. 149]:
@@
  6. .6
    Properties of the matrix $\mathbf{M}_{S}^{\bm{\Sigma}}$. Following the notation in [10], let us define the matrix $\mathbf{D}_{S}^{\bm{\Sigma}}$ as the matrix derivative of the scale function $S$ defined in (36):
@@
    The proof of these properties can be found in [10, Lemma 4.2]. Let us now focus on the property P4 to add some insight on the image of $\mathbf{M}_{S}^{\bm{\Sigma}}$. From the Property P3, we have that the dimension of $\text{Im}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$, i.e. the number of the linearly independent rows of $\mathbf{M}_{S}^{\bm{\Sigma}}$ is equal to m(m+1)/21m(m+1)/2-1. Moreover, as a direct consequence of the rank-nullity theorem, we have the following additional property:
@@
    Remarkably, the matrix $\mathbf{U}_{\bm{\Sigma}}$ is the same matrix defined in [22, Sect. IV.D] in a different set-up and it will be the bridge to prove some interesting equalities.
  7. .7
    Derivation of the semiparametric nuisance tangent space $\mathcal{T}_{g_{0}}$ Let us consider the non-parametric model introduced in (47) as:
    $\mathcal{P}_{g}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},g):g\in\bar{\mathcal{G}}\right\}.$
    Clearly, this model is the restriction to the RES-distribution of the general non-parametric model in (9) introduced in Sect. II. The aim of this appendix is to derive the related tangent space $\mathcal{T}_{g_{0}}$ (c.t. $\mathcal{T}_{3}$ in Sect. II). To this end, following [13, Sec. 4.2], [1, Sec. 3.1],[2, Sec. 2.2], we need to introduce the set of parametric sub-models of $\mathcal{P}_{g}$. Formally, the i-th parametric sub-model of $\mathcal{P}_{g}$ is defined as:
    $\mathcal{P}_{\tau_{\bm{\rho},i}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho},i}),\bm{\rho}\in\Upsilon_{i}\subseteq\mathbb{R}^{r_{i}}\right\},$ (162)
@@
    $\begin{split}\tau_{\bm{\rho},i}:\,&\mathcal{X}\times\Upsilon_{i}\rightarrow\bar{\mathcal{G}}\\ &\bm{\rho}\mapsto\tau_{i}(\mathbf{x},\bm{\rho}),\end{split}$ (163)
    is a known function parametrized by an unknown finite-dimensional vector $\bm{\rho}$. In particular, for every $i\in\mathbb{N}$, $\mathcal{P}_{\bm{\theta},\tau_{i}}$ in (162) is a parametric model satisfying the following three conditions [13, Sec. 4.2]:
    C0)
    $\tau_{\bm{\rho},i}:\mathcal{X}\times\Upsilon_{i}\rightarrow\bar{\mathcal{G}}$ is a smooth parametric map $\forall i\in\mathbb{N}$,
@@
    $\mathcal{P}_{\tau_{i}}\subseteq\mathcal{P}_{g},\;\forall i\in\mathbb{N}$,
@@
    $p_{0}(\mathbf{x})\in\mathcal{P}_{\tau_{i}}$, i.e. $\forall i\in\mathbb{N}$ there exists a vector $\bm{\rho}_{0}\in\Upsilon_{i}$ such that $p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho}_{0},i})=p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},g_{0})$.
    Intuitively, a parametric sub-model $\mathcal{P}_{\tau_{i}}$ can be though as a finite-dimensional approximation of the non-parametric model $\mathcal{P}_{g}$. The purpose of using a parametric sub-model lies in the fact that its tangent space is well-defined as shown in (12) as:
    $\mathcal{H}\supseteq\mathcal{T}_{\bm{\rho}_{0},i}\triangleq\mathrm{Span}\{[\mathbf{s}_{\bm{\rho}_{0},i}]_{1},\ldots,[\mathbf{s}_{\bm{\rho}_{0},i}]_{r_{i}}\},$ (164)
    where $\mathbf{s}_{\bm{\rho}_{0},i}=\nabla_{\bm{\rho}}\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho}_{0},i})$ is the score vector of $\bm{\rho}_{0}\in\Upsilon_{i}$ in the ithi-th parametric sub-model $\mathcal{P}_{\tau_{i}}$. Consequently, according to [14] and [13, Sec. 4.4], the tangent space $\mathcal{T}_{g_{0}}$ can be defined as the closure  [^4] The closure $\overline{\mathcal{A}}$ of a set $\mathcal{A}$ is defined as the smallest closed set that contains $\mathcal{A}$, or equivalently, as the set of all elements in $\mathcal{A}$ together with all the limit points of $\mathcal{A}$. of the union of all the (parametric) tangent spaces $\mathcal{P}_{\tau_{i}}$:
    $\mathcal{T}_{g_{0}}=\overline{\bigcup\nolimits_{i\in\mathbb{N}}\mathcal{T}_{\bm{\rho}_{0,i}}}\subseteq\mathcal{H}.$ (165)
    Equivalently, $\mathcal{T}_{g_{0}}\subseteq\mathcal{H}$ is the subspace  [^5] The closure of a union of linear spaces doesn’t need to be linear, in general. However, as discussed in [48, Assumption S] and [13, Sec. 4.4, Remark 5]), $\mathcal{T}_{g_{0}}$ is a linear space in the vast majority of the non-pathological statistical models. of $\mathcal{H}$ composed by all the functions $h\in\mathcal{T}_{g_{0}}\subseteq\mathcal{H}$ for which there exists a sequence $\{\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i},\mathbf{c}_{i}\in\mathbb{R}^{r_{i}}\}_{i\in\mathbb{N}}$ such that $h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}}||^{2}=E_{0}\left\{(h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}})^{2}\right\}\rightarrow 0$.
    Now that we have the theoretical and formal framework, let us go back to the application at hand. Specifically, we have to show that $\mathcal{T}_{g_{0}}$ can actually be expressed as in (75):
    $\begin{split}\mathcal{T}_{g_{0}}=\left\{h\in\mathcal{H}|h(\mathbf{x})=h(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})),a.s.\;\mathbf{x}\in\mathcal{X}\right\},\end{split}$ (166)
    where, from (32), $Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})=s_{0}^{-1}(\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{V}_{S,0}^{-1}(\mathbf{x}-\bm{\mu}_{0})$.
@@
    The proof follows from some small modifications of the one in [13, Theo. 4.4]. Specifically, we need to show that:
@@
      Any element of $\mathcal{T}_{\bm{\rho}_{0},i},\forall i\in\mathbb{N}$ is an element of $\mathcal{T}_{g_{0}}$, i.e. $\mathcal{T}_{\bm{\rho}_{0},i}\subset\mathcal{T}_{g_{0}},\forall i\in\mathbb{N}$,

### v3

@@ VIII Conclusion
The semiparametric statistical efficiency in estimation problems for elliptically symmetric distributed data was analyzed in this paper. In particular, we studied the impact of finite and infinite-dimensional nuisance parameters can have on the estimation of the parameters of interest which, in the case of elliptical distributions, are contained in the location vector $\bm{\mu}$ and/or in the covariance matrix $\bm{\Sigma}$. The profound and counter-intuitive result that emerged is that, in the presence of specific finite-dimensional nuisance parameters, semiparametric efficiency can be equivalent to parametric efficiency. Specifically, in the case of elliptical distributions, not knowing the density generator does not cause any loss of efficiency when estimating $\bm{\mu}$ or a scaled version of $\bm{\Sigma}$. This result had already been demonstrated by Hallin and Paindaveine in [10, 11] by leveraging an asymptotic approximation of the projection operator as a conditional expectation with respect to an appropriate invariant sub-$\sigma$ algebra. Unlike these works, in this paper we provide an explicit derivation of such projection operator. Furthermore, as an advancement over the state of the art, we analyzed the case, important in many applications, in which the parameters of interest and the finite-dimensional nuisance parameters are given by a parametrization of the location vector and of the covariance matrix. A general condition that the parameterization in question must satisfy in order for the semiparametric efficiency to be equal to the parametric efficiency has been derived in this paper. This condition therefore allows us to test this property for any particular parameterizations. Two examples were investigated here, including the well-known low-rank parameterization, often arising in many practical signal processing applications. The paper concluded with a section in which the results derived for RES distributions are extended to the case of C-CES and NC-CES distributions. The natural follow-up to this paper will be on the development of semiparametric estimators capable of achieving parametric efficiency. A promising approach for achieving this goal is that of rank-based ($R$-) estimators [28, 30, 50].
Figure 1: MSE indices and bounds for $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$ vs $\nu$ ($n=100$).
Figure 2: MSE indices and bounds for $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ vs $\nu$ ($n=100$).
Figure 3: MSE indices and bounds for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ vs $\nu$ ($n=100$).
@@
    For a given $\mathbf{h}\in\mathcal{H}_{q}$, the projection $\Pi(\mathbf{h}|\mathcal{H}_{q})$ is defined as the vector in $\mathcal{H}_{q}$ such that [18, Theo. 5.1]:
@@
    The $q$-replicating Pythagorean theorem [15, Theo. 3.3]: Let $\mathcal{H}_{q}$ and $\mathcal{U}_{q}\subseteq\mathcal{H}_{q}$ the $q$-replicating Hilbert space and subspace defined as in Lemma .1. Let us take $\mathbf{h}\in\mathcal{H}_{q}$ and $\mathbf{u}\in\mathcal{U}_{q}$. If $\mathbf{h}\perp\mathbf{u}$, then:
@@
    Proof of Lemma II.2. Let us start by noticing that the closed subspaces $(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}$ and $\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}$ are orthogonal. In fact, it is immediate to verify that, each element $h\in\mathcal{T}_{2}+\mathcal{T}_{3}$ can be written as $h=h_{2}+h_{3}$ with $h_{2}\in\mathcal{T}_{2}$ and $h_{3}\in\mathcal{T}_{3}$. Moreover, $\forall g\in\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}$ we have that $g\in\mathcal{T}_{2}^{\perp}$ and $g\in\mathcal{T}_{3}^{\perp}$ and consequently $h\perp g$. In addition, $\mathcal{T}_{2}^{\perp}$ can be expressed as the (direct) sum of the and these two closed subspaces, i.e. $\mathcal{T}_{2}^{\perp}=(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}+(\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp})$. Consequently, from the property (20), we get:
@@
    Projection onto two subspaces $\mathcal{D}+\mathcal{T}$, where $\mathcal{D}$ has finite dimension. The following theorem has been proposed in [1, Theo. 5, A.4]. Let (,,)($\mathcal{H}$$\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$) and Hilbert space. Let $\mathcal{T}\subset\mathcal{H}$ be an infinite-dimensional subspace while $\mathcal{H}\supset\mathcal{D}=\mathrm{Span}\{s_{1},\ldots,s_{d}\}$ is a finite $d$-dimensional subspace. Then, the orthogonal projection of an element $h\in\mathcal{H}$ onto $\mathcal{D}+\mathcal{T}$ can be expressed as:
    $\Pi(h|\mathcal{D}+\mathcal{T})=\Pi(h|\mathcal{T})+\Pi(h|\mathrm{Span}\{r_{1},\ldots,r_{k}\}),$ (157)
    where $r_{i}\triangleq s_{i}-\Pi(s_{i}|\mathcal{T})$ and $k\triangleq\mathrm{dim}(\mathrm{Span}\{r_{1},..,r_{k}\})\leq d$. Moreover $k=d$ if and only if $\mathcal{D}\cap\mathcal{T}=\{0\}$.
    Proof:
    To prove this expression, let us start by noticing that any element in $\{s_{i}\}_{i=1}^{d}$ can be decomposed as:
    $s_{i}=[\Pi(s_{i}|\mathcal{T})]+[s_{i}-\Pi(s_{i}|\mathcal{T})]\triangleq t_{i}+r_{i},$ (158)
    where, by definition, $t_{i}\in\mathcal{T}$ and $r_{i}\in\mathcal{T}^{\perp}$, $i=1,\ldots,d$. As a consequence, any element $l\in\mathcal{D}$ can be decomposed as:
    $l=\sum\nolimits_{i=1}^{d}\alpha_{i}t_{i}+\sum\nolimits_{i=1}^{d}\alpha_{i}r_{i},\;\alpha_{i}\in\mathbb{R},$ (159)
    where the first term belongs to $\mathcal{T}$ while the second belongs to $\mathrm{Span}\{r_{1},\ldots,r_{d}\}\subset\mathcal{T}^{\perp}$ and then:
    $\mathcal{D}+\mathcal{T}=\mathrm{Span}\{r_{1},\ldots,r_{d}\}+\mathcal{T}.$ (160)
    In general, the elements $\{r_{1},\ldots,r_{d}\}$ may be linearly dependent, so we can extract a basis $\{r_{i_{1}},\ldots,r_{i_{k}}\}$ such that $\mathrm{Span}\{r_{1},\ldots,r_{d}\}=\mathrm{Span}\{r_{i_{1}},\ldots,r_{i_{k}}\}$. For the sake of simplicity and without loss of generality, let us suppose the $k$ linearly independent elements are the first ones, then we have:
    $\mathcal{D}+\mathcal{T}=\mathrm{Span}\{r_{1},\ldots,r_{k}\}+\mathcal{T}.$ (161)
    Finally, according to [18, eq. (5.7)] or [15, Sect. 2.4, Ex. 1], the following explicit expression holds:
    $\Pi(h|\mathcal{D}+\mathcal{T})=\Pi(h|\mathcal{T})+\Pi(h|\mathrm{Span}\{r_{1},\ldots,r_{k}\})=\Pi(h|\mathcal{T})+\mathbf{p}^{T}\mathbf{G}^{-1}\mathbf{r},$ (162)
    where $\mathbf{p}\triangleq(p_{1},\ldots,p_{k})^{T}$ with $p_{i}\triangleq\left\langle h,r_{i}\right\rangle_{\mathcal{H}}$, $\mathbf{G}]_{ij}\triangleq\left\langle r_{i},r_{j}\right\rangle_{\mathcal{H}}$, for $i=1,\ldots,k$ and $\mathbf{r}\triangleq(r_{1},\ldots,r_{k})^{T}$. Let us now focus on finding a geometric (necessary and sufficient) condition to have $k=d$, that is verified if and only if $\{r_{1},\ldots,r_{d}\}$ are linearly independents. By definition, we have that:
    $\mathrm{Span}\{r_{1},\ldots,r_{d}\}=\Pi(\mathcal{D}|\mathcal{T}^{\perp}).$ (163)
    Specifically, we can define a linear map $P$ as:
    $\begin{split}P:\;&\mathcal{D}\rightarrow\mathcal{T}^{\perp}\\ &l\mapsto\Pi(l|\mathcal{T}^{\perp}).\end{split}$ (164)
    Therefore $k=\mathrm{dim}(\mathrm{Span}\{r_{1},\ldots,r_{d}\})=\mathrm{dim}(\mathrm{Im}\{P\})$. Form the rank-nullity theorem, we also have that $\mathrm{dim}(\mathrm{Im}\{P\})+\mathrm{dim}(\mathrm{ker}\{P\})=d$. Consequently:
    $k=d\Leftrightarrow\mathrm{dim}(\mathrm{ker}\{P\})=0.$ (165)
    Since, by definition $\mathrm{ker}\{P\}=\{l\in\mathcal{D}|\Pi(l|\mathcal{T}^{\perp})=0\}=\{l\in\mathcal{D}|\Pi(l|\mathcal{T})=l\}$, we immediately have that $\mathrm{ker}\{P\}=\mathcal{D}\cap\mathcal{T}$ and consequently:
    $k=d\Leftrightarrow\mathcal{D}\cap\mathcal{T}=\{0\}.$ (166)
  5. .5
    Implicit derivation of $\bm{\nabla}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}$. The following calculation has been derived in [11, Sect. 4]. Let us start by defining the mapping $v_{11}^{S}:\mathbb{R}^{m(m+1)/2-1}\mapsto\ \mathbb{R}$ implicitly defined by the constraint $S(\mathbf{V}_{S})\equiv S(v_{11}^{S}(\underline{\mathrm{vecs}}(\mathbf{V}_{S})),\underline{\mathrm{vecs}}(\mathbf{V}_{S}))=1$. Thanks to the implicit function theorem, under Assumptions A1, A2, A3, this mapping exists, is unique and continuously differentiable around a given $\mathrm{vecs}(\mathbf{V}_{S})$. Then, we can differentiate both side of $S(v_{11}^{S}(\underline{\mathrm{vecs}}(\mathbf{V}_{S})),\underline{\mathrm{vecs}}(\mathbf{V}_{S}))=1$ w.r.t. $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$ to get:
@@
    The scale $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ implies that $|\mathbf{V}_{S}|=1$. Taking as before the differential on both side, we have [25, p. 149]:
@@
  7. .7
    Properties of the matrix $\mathbf{M}_{S}^{\bm{\Sigma}}$. The following properties have been derived in [11, 10]. Let us define the matrix $\mathbf{D}_{S}^{\bm{\Sigma}}$ as the matrix derivative of the scale function $S$ defined in (36):
@@
    The proof of these properties can be found in [11, Lemma 4.2]. Let us now focus on the property P4 to add some insight on the image of $\mathbf{M}_{S}^{\bm{\Sigma}}$. From the Property P3, we have that the dimension of $\text{Im}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$, i.e. the number of the linearly independent rows of $\mathbf{M}_{S}^{\bm{\Sigma}}$ is equal to m(m+1)/21m(m+1)/2-1. Moreover, as a direct consequence of the rank-nullity theorem, we have the following additional property:
@@
  8. .8
    Derivation of the nuisance tangent spaces $\mathcal{T}^{u}_{g_{0}}$ and $\mathcal{T}^{c}_{\bar{g}_{0}}$. Let us start by considering the unconstrained non-parametric sub-model $\mathcal{P}^{u}_{g}$ given in (49) and derive the relevant tangent space $\mathcal{T}^{u}_{g_{0}}$. To this end, following [15, Sec. 4.2], [1, Sec. 3.1],[2, Sec. 2.2], we need to introduce a set of parametric sub-models of $\mathcal{P}^{u}_{g}$. Formally, an $i$-th parametric sub-model of $\mathcal{P}^{u}_{g}$ is defined as:
    $\mathcal{P}^{u}_{\tau_{i,\bm{\rho}}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},\tau_{i,\bm{\rho}}),\bm{\rho}\in\Upsilon_{i}\subseteq\mathbb{R}^{r_{i}}\right\},$ (180)
@@
    $\begin{split}\tau_{i,\bm{\rho}}:\,&\mathcal{X}\times\Upsilon_{i}\rightarrow\mathcal{G}\\ &\bm{\rho}\mapsto\tau_{i}(\mathbf{x},\bm{\rho}),\end{split}$ (181)
    is a known function parametrized by an (artificial) unknown finite-dimensional vector $\bm{\rho}$. In particular, for every smooth parametric map $\tau_{i}(\cdot,\bm{\rho}):\Upsilon_{i}\rightarrow\mathcal{G}$, $\mathcal{\mathcal{P}}^{u}_{\tau_{i,\bm{\rho}}}$ in (180) is a parametric model satisfying the following two conditions [15, Sec. 4.2]:
@@
    $\mathcal{P}^{u}_{\tau_{i,\bm{\rho}}}\subseteq\mathcal{P}^{u}_{g},\;\forall i\in\mathbb{N}$,
@@
    $p_{0}(\mathbf{x})\in\mathcal{P}^{u}_{\tau_{i,\bm{\rho}}}$, i.e. $\forall i\in\mathbb{N}$ there exists a vector $\bm{\rho}_{0}\in\Upsilon_{i}$ such that $\color[rgb]{0,0,0}p_{X}(\mathbf{x}|\bm{\gamma}_{0},\tau_{i,\bm{\rho}_{0}})=p_{X}(\mathbf{x}|\bm{\gamma}_{0},g_{0})\triangleq p_{0}(\mathbf{x})$}.
    Intuitively, a parametric sub-model $\mathcal{P}^{u}_{\tau_{i,\bm{\rho}}}$ can be thought as a finite-dimensional approximation of the non-parametric model $\mathcal{P}^{u}_{g}$. The purpose of using a parametric sub-model lies in the fact that its tangent space is well-defined as shown in (12) as:
    $\mathcal{H}\supseteq\mathcal{T}^{u}_{i,\bm{\rho}_{0}}\triangleq\mathrm{Span}\{[\mathbf{s}_{\bm{\rho}_{0},i}]_{1},\ldots,[\mathbf{s}_{\bm{\rho}_{0},i}]_{r_{i}}\},$ (182)
    where $\mathbf{s}_{\bm{\rho}_{0},i}={\color[rgb]{0,0,0}\nabla_{\bm{\rho}}\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},\tau_{i,\bm{\rho}_{0}})}$ is the score vector of $\bm{\rho}_{0}\in\Upsilon_{i}$ in the $i$-th parametric sub-model $\mathcal{P}_{\tau_{i}}^{u}$. Consequently, according to [16] and [15, Sec. 4.4], the tangent space $\mathcal{T}^{u}_{g_{0}}$ can be defined as the closure in $\mathcal{H}$  [^9] The closure $\overline{\mathcal{A}}$ of a set $\mathcal{A}$ is defined as the smallest closed set that contains $\mathcal{A}$, or equivalently, as the set of all elements in $\mathcal{A}$ together with all the limit points of $\mathcal{A}$. of the union of all the (parametric) tangent spaces $\mathcal{T}^{u}_{i,\bm{\rho}_{0}}$:
    $\mathcal{T}^{u}_{g_{0}}=\overline{\bigcup\nolimits_{i\in\mathbb{N}}\mathcal{T}^{u}_{i,\bm{\rho}_{0}}}\subseteq\mathcal{H}.$ (183)
    Equivalently, $\mathcal{T}^{u}_{g_{0}}\subseteq\mathcal{H}$ is the subspace  [^10] The closure of a linear spaces doesn’t need to be linear, in general. However, as discussed in [56, Assumption S] and [15, Sec. 4.4, Remark 5]), $\mathcal{T}^{u}_{g_{0}}$ is a linear space in the vast majority of the non-pathological statistical models. of $\mathcal{H}$ composed by all the functions $\color[rgb]{0,0,0}h\in\mathcal{H}$} for which there exists a sequence $\{\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i},\mathbf{c}_{i}\in\mathbb{R}^{r_{i}}\}_{i\in\mathbb{N}}$ such that $h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}}||^{2}=E_{0}\left\{(h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}})^{2}\right\}\rightarrow 0$.
    Now that we have the theoretical and formal framework, let us go back to the application at hand. Specifically, we have to show that $\mathcal{T}^{u}_{g_{0}}$ can actually be expressed as in (77):
    $\begin{split}\mathcal{T}^{u}_{g_{0}}&=\left\{h\in\mathcal{H}|h(\mathbf{x})=h((\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{V}_{S,0}^{-1}(\mathbf{x}-\bm{\mu}_{0})),a.s.\;\mathbf{x}\in\mathcal{X}\right\}\\ &=\left\{h\in\mathcal{H}|h(\mathbf{x})=h(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x})),a.s.\;\mathbf{x}\in\mathcal{X}\right\}\\ &=\left\{h\in\mathcal{H}|h\;\text{is $\sigma(\mathcal{Q})$-measurable}\right\}.\end{split}$ (184)
@@
    The proof follows from some small modifications of the one in [15, Theo. 4.4]. Specifically, we need to show that:
@@
      Any element of $\mathcal{T}^{u}_{i,\bm{\rho}_{0}},\forall i\in\mathbb{N}$ is an element of $\mathcal{T}^{u}_{g_{0}}$, i.e. $\mathcal{T}^{u}_{i,\bm{\rho}_{0}}\subset\mathcal{T}^{u}_{g_{0}},\forall i\in\mathbb{N}$,